OpenStudy (anonymous):

Let g(t)=(t+3)ln(t+5) Find the derivative, g'(t) , and the second derivative, g"(t) . Is my first derivative right? g'(t)=ln(t+5)+(1/ln(10))+(3/tln(10))

OpenStudy (anonymous):

Incorrect... The derivative of ln(t+5) is just 1/(t+5) G'(t) = ln(t+5) + (t+3)/(t+5) Product Rule is used.

OpenStudy (anonymous):

Oh okay so would the second derivative would be g"(t)=(1/t ln(10))+(2/(t+5)^2)?